Literaturauswahl

zur Vorlesung Elementare Algorithmen in der Algebra SS 2000

• W. W. Adams, P. Loustaunau, An introduction to Gröbner bases. AMS graduate studies in math. 3 (1994)

• Eric Bach, Jeffrey Shallit, Algorithmic number theory. Cambridge (1996)

• T. Becker, V. Weispfennig, Gröbner bases. Springer (1993)

• Johannes Buchmann, Einführung in die Kryptographie. Springer (1999)

• P.M. Cohn, Algebra. Volume 1, Wiley (1995)

• Henri Cohen, A course in computational algebraic number theory. Springer Graduate Text in Mathematics 138 (1993)

• N. Koblitz, Algebraic aspects of Cryptography, Springer (1997)

• Michael Pohst und Hans Zassenhaus, Algorithmic algebraic number theory. Cambridge university press (1989) (paperback, 1997).

• Hans Riesel, Prime numbers and computer methods of factorization. Birkhaeuser (1994)

• Fuzhen Zhang, Matrix theory. Springer (1999)